answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Is standard deviation an absolute value?

No. Standard deviation is not an absolute value.

The standard deviation is often written as a single positive value (magnitude), but it is really a binomial, and it equals both the positive and negative of the given magnitude.

For example, if you are told that for a population the SD is 5.0, it really means +5.0 and -5.0 from the population mean. It defines a region within the distribution, starting at the lower magnitude (-5.0) increasing to zero (the mean), and another region starting at zero (the mean) and increasing up to the upper magnitude (+5.0). Both regions together define the (continuous) region of standard deviation from the mean value.

What is the definition of standard deviation and variance?

The set of X1, X2, ..., XN is called X. Given that mean(X), is the sum of all X divided by N, the variance of X is mean((Xi - mean(X))2).

The standard deviation of X is the square root of the variance.

Company buys calculators from a Korean supplier The probability of a defective calculator is 10 percent If three calculators are selected at random what is the probability that none of the calculato?

The question has been truncated.

I suspect it either should end: "...none of the calculators work," in which case:

for all 3 to be faulty the probability is 1/10 × 1/10 × 1/10 = 1/1000 = 0.001 = 0.1 %

or: "...none of the calculators are faulty," in which case:

for all 3 to not be faulty the probability = 9/10 × 9/10 × 9/10 = 729/1000 = 0.729 = 72.9 %

What is the probability of rolling a sum of 11 with 2 dice?

The probability of rolling a sum of 11 with 2 dice is: P(11) = 1/18.

For explanation see answer to question: "What is the probability of rolling 7 or 11 with 2 dice?".

Read more: What_is_the_probability_of_rolling_7_or_11_with_2_dice

What is certain probability?

If something is certain, then the probability of that something is 1.

Suppose you roll two dice Find the probability of rolling a sum of 5?

Assuming these are six sided dice, there should be 4 ways to roll a sum of 5.

1 + 4 = 5

4 + 1 = 5

2 + 3 = 5

3 + 2 = 5

There are 36 (6^2) possible combinations of rolling two six sided dice.

That should mean 4 out of 36 rolls should give us a sum of 5 or reduced to 1/9.

Suppose you roll two dice Find the probability of rolling a sum of 6?

Find the possible combinations of rolling a six with two dice:

1 + 5 = 6

5 + 1 = 6

2 + 4 = 6

4 + 2 = 6

3 + 3 = 6

3 + 3 = 6

Six in total.

Find the number of possible combinations;

Number of Sides on the Dice (6) to power of how many dice (2) = 36

Number of Times the Dice will sum to Six (6 times)

--------------------------------------------------

Number of Possible outcomes

This will give us 6 out of 36 rolls.

Or reduced 1/6 rolls

If you draw a card from a standard deck of cards what is the probability of drawing a red card?

There are 2 red suits and 2 black suits. Therefore the probability of drawing a red card is 1/2. Or 50% chance.

What is the square of the standard deviation called?

The square of the standard deviation is called the variance. That is because the standard deviation is defined as the square root of the variance.

What is the probability of getting at most one head when two coins are tossed?

Sample space for two coins tossed is:

HH

HT

TH

TT

Therefore at most one head is HT TH TT or 3/4 or 0.75.

What is the probability of rolling a 3 and then 1 of six side dice?

The probability of rolling the 3 is (1/6).

The probability of rolling the 1 is (1/6).

The probability of rolling the 3 and then the 1 is (1/6) x (1/6) = (1/36) = about 2.78% (rounded)

If you toss a coin 50 times what will it land on?

Probably on the floor at least once.

It's possible for the results to be anything between 50 heads and 50 tails.

The most probable result ... if the coin is honest, fair, and balanced, like Fox News #@%$&#*&@ ...

is 25 heads and 25 tails.

There is no way to predict exactly what the result will be. If it were possible to know in advance,

then coin-tosses could not be used to gamble or to decide the answer to a question.

Probability of having a boy or a girl?

The probability of having a boy or a girl is always 50/50 because the female has to X sex chromosomes and the male has one X and one Y sex chromosome. If you do a Punnet Square, there will be two squares with XY and two squares with XX. With four squares in all that makes a 50/50 chance.

If you toss 3 coins what are the odds of getting exactly 3 heads?

The odds of getting 3 heads in a toss of 3 coins is 1 in 8, or 0.125.

Each coin is probabalistically unrelated to each other, so you simply multiply the odds for each coin. 1 in 2 times 1 in 2 times 1 in 2 is 1 in 8.

What are the rules of probability?

Basic Rules of Probability:

1) The probability of an event (E) is a number (fraction or decimal) between and including 0 and 1. (0≤P(E)≤1)

2) If an event (E) cannot occur its probability is 0.

3) If an event (E) is certain to occur, then the probability if E is 1. This means that there is a 100% chance that something will occur.

4) The sum of probabilities of all the outcomes in the sample space is 1.

Addition Rules/Formulas:

When two events (A and B) are mutually exclusive, meaning that they can't occur at the same time or they have no outcomes in common, the probability that A or B will occur is:

P(A or B)= P(A)+P(B)

If A and B are not mutually exclusive, then:

P(A or B)= P(A)+P(B)-P(A and B)

Multiplication Rules/Formulas:

When two events (A and B) are independent events, meaning the fact that A occurs does not affect the probability of B occurring (for example flipping a coin, rolling a die, or picking a card), the probability of both occurring is:

P(A and B)= P(A)P(B)

Conditional Probability-When two events are dependent (not independent), the probability of both occurring is:

P(A or B)= P(A)P(B|A)

Note: P(B|A) does not mean B divided by A but the probability of B after A.

How many different arrangements can be made with the letters in the word Iowa?

There are 4 letters in IOWA, all non repeating. Arrangements are 4! or 24.

When an event in certain to occur What is its probability?

p = 1

For example in popular language:

An event certain to occur is that the sun will rise each day. The probability of the sun rising is 100%.

But, as you are well aware, it's not the sun rising but the earth turning and it's only going to happen for another 10 billion years.

What is irregular variation?

an irregular variation is one which occurs predictably and at random, usually over a very short time period. it is impossible to allow for irregular variations in forecasting rather, they are used to explain abnormalities after the event.

hope this helps :-)